Write the statement as an equation. (Use k as the constant of variation.)
W varies jointly with r and the square of s.
W=???

Answer by Theo(13342)

w varies jointly with r and the square of s.

the formula for joint variation is z = k * x * y

let w = z
let r = x
let s^2 = y

you get:

w = k * r * s^2

here's a reference.

http://www.mathwords.com/jk/joint_variation.htm